OWG Meeting 29-Jun-2016

From GlueXWiki
Jump to: navigation, search

Location and Time

Room: CC F326

Time: 2:00pm-3:00pm


You can connect using BlueJeans Video conferencing (ID: 120 390 084). (Click "Expand" to the right for details -->):

(if problems, call phone in conference room: 757-269-6460)

  1. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 120390084.
  2. To join via a Web Browser, go to the page https://bluejeans.com/120390084.
  3. To join via phone, use one of the following numbers and the Conference ID: 120390084
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. More information on connecting to bluejeans is available.

Previous Meeting


  1. Announcements
    • VME crates powered on (but not detectors)
    • 8 farm nodes ordered
    • 36-port IB switch (56GB/s) w/ cables : "Buyer Assigned"
    • RAID disk not ordered
    • gluon43 sent for repair
    • Future OS update to RHEL6.5 - No update
    • /gluex disk filled
  2. DAQ
    • pedestal check vs. testing mode
    • CODA 3.0.6
  3. Front-end Firmware Status
  4. L3 Status (meetings)
  5. AOT

Recharge Wednesday: Ice cream novelties


Attendees: David L., Simon T., Sergey F., Bryan M., Beni Z., Dave A., Vardan G., Naomi J., Sean D., William G., Carl T.


  • /gluex disk filled earlier in week due to log files from FCAL. These have been growing since the FCAL was turned off due to error messages
    • Removing log files doesn't immediately free disk space due to snapshot maintaining hard links
    • Paul L. added some space to partition to get us through next several days until snapshot links are deleted
    • Problem may reoccur if someone doesn't remember to watch log files
    • May need more robust system for managing logs to ensure they don't overwrite disk. (David will discuss with Hovanes et al.)


  • Crates are on, but new feature in ROL prevents DAQ from starting if pedestals in trigger related crates are not correct.
    • With HV/LV off, the pedestals in BCAL are not correct so error flag is set to prevent DAQ from starting
    • Sergey spoke with Alex S. and Alex will implement a configuration parameter that allows the feature to be disabled for special testing
    • In meantime, Sergey will disable in ROL for DAQ testing that doesn't require normal L1 trigger to be working properly

Front-end Firmware Status

  • Some crates have had firmware updated (e.g. 2 fADC250 crates) but not all
    • New firmware requires ROL be linked with corresponding library. Currently all ROCs use same ROL so using mixed firmware versions is ugly to implement.
    • Will ask Alex S. about accelerating schedule to get all crates updated so we can do full set of crates to test CODA 3.06
    • Some interest in getting new cosmic data

L3 Status

  • Dave shared some results presented at last L3 meeting. Specifically, that there appears to be an optimal configuration of parsing and processing threads when doing "parsing only"

AOT: Trigger Monitoring

  • Beni noted that the trigger monitor program he put together during the last run was limited in its ability to calculate trigger rates for pulser events that happen rarely since his program only sees a fraction of events
  • Several suggestions made on how this could be improved
    • Set up blocking ET station on monitoring ET system to force more (all?) events through trigger monitor
    • TS scalers already in sync events. Set up station to get only sync events
    • Send TS scalers to Jflux system and make them available via web browser.
  • Beni will consider these options